A sample of an ideal gas has a volume of 3.60 L at 13.40 degrees Celsius and 1.30 atm. What is the volume of gas at 24.60 degrees Celsius and 0.990 atm
1
Expert's answer
2017-10-27T13:53:06-0400
Combined gas Law
P1V1 / T1 = P2V2 / T2
P1 = 1.30 atm
V1 = 3.60 L
T1 (must be in Kelvin = 13.40 deg C + 273.15) = 286.55 K
P2 = 0.990 atm
V2 = ? L
T2 = 297.75 K
Solve for V2
V2 = P1V1T2 / (P2T1) = 1.30 atm x 3.60 L x 297.75 K / (0.990 atm x 286.55 K) = 4.91 L